OpenStudy (anonymous):

Given equation A as 2x + 1/4y =3 and equation B as 2/3 x - y = 6 , which expression will give the value of y?

OpenStudy (anonymous):

u hv to use both equations

OpenStudy (mathmate):

You are looking for an expression which will eliminate x, and from which y can be solved. Note that the term 1/4y is ambiguous, which could mean (1/4)y or 1/(4y). From the context, I take it as (1/4)y. Unambiguous ways to write it would be (1/4)y, or y/4, and at least 1/4 y, as you have done so in equation B.

OpenStudy (mathmate):

The expression would be in terms of A and B. If I interpreted correctly, the equations are: \[A: 2x + \frac{1}{4}y =3 \\ and\\ B: \frac{2}{3}x - y = 6 \] we can eliminate y by adding together 4 times equation A and equation B to get \[4A: 8x + y =12\\ B: \frac{2}{3}x - y = 6\] so \[4A+B: 8\frac{2}{3}x = 18 \] from which x can be readily found. Thus the expression which will give x is 4A+B. Now it's your turn to find the expression which will give y.

OpenStudy (anonymous):

none of those are a choice A. a-3b B. a- 3/2 b C. 2a-3b D. 1/3 a+b

OpenStudy (anonymous):

D seems to be the closest though

OpenStudy (mathmate):

What I did is to find the expressions that gives x. You need to follow the same steps to find the one that gives y, so 4A + B should not be in your choices. I see a correct answer among the choices. Give it a try to work it out. You need to eliminate x to find y.